A boy is twice as old as his brother - 4 years ago he was 3 times as old as his brother. What are their present ages?
Answer by jorel1380(2518)

If boy X is twice as old as his brother Y right now, then X = 2Y. Four years ago he was three times as old as his brother, so X - 4=3(Y - 4).
X-4 = 3(Y-4)
X-4 = 3Y-12
If X=2Y, then
2Y-4=3Y-12
2Y+8=3Y
8=Y
The brother is now 8 years old, and the boy is 2(8) or 16 years old.